# Deformation dynamics of an oil droplet into a crescent shape during the intermittent motion

### v1

@@ Abstract
A paraffin droplet containing camphor and oil red O dye floating on the water surface exhibits spontaneous motion with deformation by generating a surface tension gradient around the droplet. We focused on the intermittent motion with a pronounced deformation into a crescent shape observed at specific concentrations of camphor and oil red O. We quantitatively analyzed the time changes in the droplet deformation and investigated the role of the oil red O by measuring the time-dependent oil/water interfacial tension with the pendant drop method. The observed effect can be explained by adsorption of the oil red O molecules at the oil/water interface: the interfacial tension decreases gradually after the interface formation, resulting in the dynamic deformation of the droplet. The combination of the decrease in interfacial tension and the decrease in driving force related to camphor outflow generates intermittent motion with dynamic deformation into a crescent shape.
@@ I Introduction
Self-propelled motion can be observed in many nonequilibrium systems, including living organisms. There are also nonliving systems in which the free energy is converted into mechanical motion. They are much less complex than living organisms, and the mechanism of their self-propulsion can be explained with mathematical models through physics and chemistry. Consequently, studies on nonliving self-propelled motion are crucial for deep understanding of the phenomenon on the grounds of nonequilibrium science.
@@
A camphor disk placed on the water surface is one of the well-known examples of self-propelled particles[1, 2]. After a disk is placed on the water surface, a layer of camphor molecules is formed, reducing the surface tension. Fluctuations in the concentration gradient can appear around the disk, leading to the surface tension gradient that drives the disk toward the slightly lower concentration area. Once the disk starts moving, its motion is maintained since the surface tension is lower at the rear of the disk[3, 4, 5]. Thus, a circular camphor disk moves by spontaneously breaking symmetry[6, 7]. In typical laboratory conditions camphor molecules evaporate quickly; hence, the water surface tension recovers in a short time. Therefore, the stationary motion of a camphor disk can be observed for a long time exceeding one hour. The self-propulsion of a camphor disk has been studied for more than a century[8, 9]. While fluctuations govern the direction of motion of a circular camphor disk, the motion of an asymmetric disk is determined by its shape[10, 11, 12, 13, 14, 15, 16, 17]. For example, it has been shown both theoretically and experimentally that an elliptical object tends to move in the direction of its minor axis[16, 17]. There are also several systems in which the particle shape determines the motion[18, 19, 20, 21].
@@
For the self-propelled motion of a solid object, its shape determines the motion, but the shape is not affected by the motion since it is fixed. On the contrary, the feedback between motion and shape can be anticipated for self-propelled soft objects like droplets because motion can modify the object shape, and the shape affects its motion. The interactions between motion and shape of self-propelled droplets were studied in a number of papers[22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. The relationship between shape and motion was also observed in biological systems. For example, a keratocyte cell moves with deformation, and it is known experimentally and theoretically that the direction of motion strongly correlates with its shape[35, 36]. Therefore, the mechanism of the coupling between motion and deformation is an important problem in self-propulsion in non-equilibrium systems. There have been several theoretical reports in which a generic model was constructed for the coupling between motion and deformation by considering the symmetric properties[37, 38, 39]. However, these studies are concerned with small magnitudes of deformation from a circle. As far as the authors know, there has been no universal description from the theoretical aspect that can be applied to large deformation.
@@
It has been recently reported that a paraffin droplet with camphor floating on the water surface moves spontaneously. Such a droplet shows complex behavior that can be controlled by the concentration of the oil red O dye in the paraffin[40, 41]. For moderate oil red O concentrations, a droplet moves spontaneously with dynamic deformation, whose shape cannot be described by a perturbation from a circular shape. A quantitative analysis of the phenomenon should help us to better understand the self-propulsion for systems showing big shape deformation.
@@
In the presented study, we investigated the mechanism of deformation of a paraffin droplet containing camphor and oil red O, in which intermittent motion with dynamic deformation can be observed depending on the concentrations of camphor and oil red O and the elapsed time from putting the droplet on the water surface. We mainly focused on the change between the crescent and circular shapes observed during the intermittent motion. We introduced the quantities that allow us to measure the degree of deformation and evaluated them by image processing of the experimental results. Then, we discussed the relationship between these quantities and the droplet speed. We also measured the interfacial tension between water and oil containing oil red O and proposed the deformation mechanism towards a crescent shape coupled with intermittent motion.
@@ II.1 Spontaneous motion of a droplet
(+)-Camphor and oil red O (dye) were purchased from FUJIFILM Wako Pure Chemical Corporation (Japan), and paraffin oil was purchased from Sigma-Aldrich (USA). The chemicals were used without further purification. Ion-free water was prepared with Elix UV3 (Merck, Germany). We prepared a paraffin solution of camphor and oil red O at different concentrations by stirring using a vortex mixer and sonication.
We observed the time evolution of a 20 $\mu$L droplet located on the water surface in a Petri dish with an inner diameter of 194.5 mm. The time evolution of the system was recorded from above for an hour at 50 fps with the CMOS camera (STC-MBS43U3V, OMRON SENTECH Co., LTD., Japan) equipped with the objective lens (3Z4S-LE SV-0814H, OMRON Co., LTD., Japan). The Petri dish was illuminated by a light panel placed below.
All experiments were performed at room temperature ($2\pm 2$ ${}^{\circ}$). Analyses of the recorded videos were performed by an image-processing software ImageJ[42].
@@ II.2 Interfacial tension between oil and water
The interfacial tension between pure water and oil was measured using the pendant drop method. Paraffin oil containing oil red O was ejected through a hook-shaped needle (4983 PD 22G, Kyowa Interface Science Co., Japan) into the aqueous phase. The droplet image was recorded from the side with the above-mentioned camera equipped with the objective lens (TEC-M55, CBC Co., Ltd., Japan). It was illuminated by the LED lamp (AS3000, AS ONE Co., Japan) located at the opposite side to the camera. The interfacial tension was calculated by fitting the shape of an oil droplet to a theoretical curve. The details are described in Appendix A.
@@ III.1 Spontaneous motion of a droplet
A 20 $\mu$L droplet of paraffin oil with dissolved camphor (concentration 0.06 M) and oil red O (concentration 1.0 g/L) exhibited three different types of motion depending on the time $t$ elapsed from the moment when the droplet was placed on the water surface. At the beginning of the experiment, the droplet showed self-propelled motion along a straight line perturbed by reflections with the dish wall, as shown in Fig. 1(a-1). Approximately at $t=30$ s, the droplet began to rotate along the wall (Fig. 1(a-2)). Then, at $t\sim 2000$ s, the droplet motion switched to intermittent motion (Fig. 1(a-3)). At this stage, short intervals of time when a droplet moved (less than $10$ s) were separated by periods when it was standing. The droplet had a circular shape when it did not move, and it became elongated and exhibited deformation into a crescent shape during intermittent motion.
@@
Figure 1: Typical long-time behavior of a paraffin droplet containing camphor (concentration 0.06 M) and oil red O (concentration 1.0 g/L) moving on a water surface. The time $t=0$ corresponds to the moment when the droplet was placed on the water surface. (a) Snapshots of the droplet motion at around $t\sim 0$ s (a-1), $t\sim 600$ s (a-2), and $t\sim 2100$ s (a-3). The superposed images of the droplet are shown with the time difference of 0.5 s (a-1) and 1.0 s (a-2, a-3). The red lines indicate the trajectories of the droplet center of mass (COM) position. (b) Droplet speed $v$ as a function of time $t$.
@@
Hereafter, we focus on the intermittent motion shown in Fig. 1(a-3). We defined a single phase of intermittent motion based on the droplet speed as a function of time (cf. Fig. 3(b)). If the droplet stopped ($v\leq 2$ mm/s) for more than 4 s, then we assumed that the droplet was at the rest state. If the droplet at the rest state was accelerated and its speed overcame 10 mm/s, then we considered this moment as the start of the intermittent phase. In the case that the time duration between the start of motion and the return to the rest state was less than 10 s, we defined that the droplet had performed an intermittent motion, and we analyzed the time series during each intermittent motion in detail. To minimize the boundary effect we did not analyze the intermittent motion if the droplet got closer than 20 mm to the wall of the Petri dish. We also excluded the data after a fission.
@@
In the following analysis, we focused on the correlations between the droplet speed and dynamic deformation into a crescent shape observed during the intermittent motion (Fig. 1(a-3)). The observed droplet elongation was always perpendicular to the direction of motion. To study the deformation dynamics quantitatively, we introduced the local coordinates associated with the direction of motion. Each frame of the observed time evolution was processed so that pixels representing the droplet were black (the logical value $1$) and all the others were white (the logical value 00). For each frame, we calculated the droplet center $\bm{r}_{\mathrm{c}}(t)$ as the center of mass (COM) of the black pixels. The set of droplet centers defined the droplet trajectory. The origin of the local coordinate system was set at $\bm{r}_{\mathrm{c}}(t)$. The $Y$-axis was defined as parallel to a fitting line to the droplet trajectory with a least-square method and directed as the net displacement vector. The $X$-axis was set perpendicular to the $Y$-axis. Using these coordinates, we defined the time-dependent aspect ratio $R=a/b$, where $b$ is the width of the droplet measured on the $Y$-axis and $a$ is the length of the droplet projected on the $X$-axis as illustrated in Fig. 2(a). The second quantity describing time-dependent droplet shape is the front-back asymmetric parameter $A=\Delta\Sigma/(2\Sigma)$, where $\Sigma$ is the area (the number of black pixels) of the droplet image and $\Delta\Sigma$ is the area of the figure obtained by applying XOR operation to the original droplet image and its image reflected with respect to the $X$-axis as shown in Fig. 2(b). It should be noted that $0\leq A\leq 1$ and $A=0$ for a droplet with mirror symmetry on the $X$-axis.
@@
Figure 3: Dynamics of a single phase of the intermittent motion of a 20 $\mu$L paraffin droplet with camphor and oil red O. The concentrations of camphor and oil red O were 0.06 M and 1.0 g/L, respectively. (a) The speed $v$ of the droplet COM, the aspect ratio $R$, and the asymmetric parameter $A$ as functions of $\Delta t$. The time $\Delta t=0$ corresponds to when $v$ has the maximum value. (b) A sequence of frames showing droplet location and deformation during the phase of intermittent motion described in (a).
@@
Figure 4: Scatter plot of $\Delta t_{A}$ and $\Delta t_{R}$, which are times for maximum values of $A(t)$ and $R(t)$ for the paraffin oil droplet with camphor and oil red O. The camphor concentration was 0.06 M and oil red O concentration was 1.0 g/L.
Next, we investigated the dependence of the droplet deformation on the oil red O concentration. We performed experiments at oil red O concentrations in the range of 0.5–1.1 g/L. The maximum values of $R(t)$ and $A(t)$ ($R_{\mathrm{max}}$ and $A_{\mathrm{max}}$), increased for oil red O concentrations larger than approximately 0.7 g/L as shown in Fig. 5.
Figure 5: (a) Maximum values of $R_{\mathrm{max}}$ and (b) $A_{\mathrm{max}}$, for $R(t)$ and $A(t)$ as functions of oil red O concentration. The camphor concentration was 0.06 M, and experiments at each oil red O concentration were repeated more than 5 times. The error bars indicate the standard deviation of the results.
@@ III.2 Interfacial tension between oil and water
To understand the role of oil red O, we measured the interfacial tension $\gamma_{\mathrm{ow}}(t)$ between the paraffin droplet containing oil red O and pure water using the pendant drop method[43] as a function of time. The measurement was performed for various oil red O concentrations. The interfacial tension gradually decreased in time after the interface was formed in the case that the oil red O concentration was higher, as shown in Fig. 6. For quantitative evaluation, we fitted time-dependent interfacial tension $\gamma_{\mathrm{ow}}(t)$ to an exponential function
@@
where $\gamma_{\mathrm{ow}}^{(0)}$ is the interfacial tension at the initial stage, $\Delta\gamma_{\mathrm{ow}}$ is the difference in interfacial tension between the initial and final stages, and $\tau$ is the characteristic time for the decrease in interfacial tension. We plotted $\gamma_{\mathrm{ow}}^{(0)}$, $\Delta\gamma_{\mathrm{ow}}$ and $\tau$ against oil red O concentration in Fig. 7. $\gamma_{\mathrm{ow}}^{(0)}$ and $\Delta\gamma_{\mathrm{ow}}$ were decreasing and increasing functions of oil red O concentration, respectively. In the case when the decrease in interfacial tension was clearly observed (oil red O concentrations: 0.5, 0.8, and 1.0 g/L), $\tau$ was approximately 800 s. In addition, we performed the same measurement for the oil droplet that contained both oil red O (1.0 g/L) and camphor (0.06 M). The decrease in interfacial tension was also observed, and $\gamma_{\mathrm{ow}}^{(0)}$ was lower compared to that for the oil droplet with oil red O (1.0 g/L) but without camphor.
Figure 6: Time evolution of the interfacial tension $\gamma_{\mathrm{ow}}(t)$ between water and paraffin oil droplet measured by the pendant drop method. The time $t=0$ corresponds to when a droplet was formed. Curve colors correspond to oil red O concentrations: 0 g/L black, 0.1 g/L yellow, 0.5 g/L orange, 0.8 g/L red, 1.0 g/L dark red. The average values of four experiments are plotted. The error bars indicate the standard deviation of the results. $\gamma_{\mathrm{ow}}(t)$ between pure water and paraffin oil droplet that contained both oil red O (1.0 g/L) and camphor (0.06 M) is plotted as blue squares.
Figure 7: Parameters fitting $\gamma_{\mathrm{ow}}(t)$ according to Eq. 1 for different concentrations of oil red O (cf. Fig. 6). (a) Interfacial tension on the initial stage $\gamma_{\mathrm{ow}}^{(0)}$. (b) Difference in interfacial tension $\Delta\gamma_{\mathrm{ow}}$ between the initial and final stages. (c) Characteristic time $\tau$ for the decrease in interfacial tension. The data for a paraffin oil droplet that contained both oil red O (1.0 g/L) and camphor (0.06 M) are shown with blue squares. The average values for four measurements are shown. The error bars indicate their standard deviations.

### v2

@@ Abstract
A paraffin droplet containing camphor and oil red O (dye) floating on the water surface shows spontaneous motion and deformation generated by the surface tension gradient around the droplet. We focused on the intermittent motion with a pronounced deformation into a crescent shape observed at specific concentrations of camphor and oil red O. We quantitatively analyzed the time changes in the droplet deformation and investigated the role of the oil red O by measuring the time-dependent paraffin/water interfacial tension with the pendant drop method. The observed effect can be explained by the active role of the oil red O molecules at the paraffin/water interface. The interfacial tension decreases gradually after the interface formation, allowing for the dynamic deformation of the droplet. The combination of the decrease in interfacial tension and the reduction in driving force related to camphor outflow generates intermittent motion with dynamic deformation into a crescent shape.
@@ I Introduction
Self-propelled motion can be observed in many nonequilibrium systems, and it is regarded as one of the attributes of life[1]. There are also nonliving systems in which the free energy is converted into mechanical motion. They are much less complex than living organisms, and the mechanism of their self-propulsion can be explained with mathematical models through physics and chemistry. Consequently, studies on nonliving self-propelled motion are crucial for a deep understanding of the phenomenon of self-motion on the grounds of nonequilibrium science.
@@
A camphor piece placed on the water surface is one of the well-known examples of self-propelled objects[2, 3]. The self-propulsion of a camphor disk has been studied for more than a century[4, 5]. After a disk is placed on the water surface, a layer of camphor molecules is formed, reducing the surface tension. Fluctuations in the camphor surface concentration can appear around the disk, leading to the surface tension gradient that drives the disk toward the lower concentration area. Once the disk starts moving, its motion is maintained since the surface tension in front of a disk is higher[6, 7, 8]. Thus, a circular camphor disk moves by spontaneous symmetry breaking[9, 10]. In typical laboratory conditions, camphor molecules evaporate quickly; hence, the water surface tension recovers in a short time. Therefore, camphor disk motion can be observed for a long time exceeding one hour. While fluctuations govern the direction of motion of a circular camphor disk, the motion of an asymmetric disk is determined by its shape[11, 12, 13, 14, 15, 16, 17, 18]. For example, it has been demonstrated both theoretically and experimentally that an elliptical object tends to move in the direction of its minor axis[17, 18]. There are also several other systems in which the particle shape determines the motion[19, 20, 21, 22].
@@
For the self-propelled motion of a solid object, its shape determines the motion. On the other hand, the feedback between motion and shape can be anticipated for self-propelled soft objects like droplets because motion can modify the object shape, and the shape affects its motion. The interactions between motion and shape of self-propelled droplets were studied in a number of papers[23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. The mechanism of the coupling between motion and deformation is an essential problem in self-propulsion in non-equilibrium systems. There have been several theoretical reports in which generic models for the coupling between motion and deformation by considering the symmetric properties were discussed[36, 37, 38]. These generic models are also applicable to the biological cell motion with deformation[39, 40]. However, these studies were concerned with small magnitudes of deformation from a circle. As far as the authors know, there has been no universal theoretical description that can be applied to large deformation.
@@
Experimental results on self-motion of a paraffin droplet with camphor floating on the water surface have been reported. Such a droplet shows complex behavior that can be controlled by the concentration of the oil red O in paraffin[41, 42]. For moderate oil red O concentrations, a droplet moves spontaneously with shape deformation that cannot be described as a small perturbation from a circular shape. A quantitative analysis of the phenomenon should help for better understanding of the self-propulsion in systems showing significant shape deformations.
@@
In the presented study, we investigated the mechanism of deformation of a paraffin droplet containing camphor and oil red O, in which intermittent motion with dynamic deformation can be observed depending on the concentrations of camphor and oil red O and the elapsed time from putting the droplet on the water surface. We mainly focused on the change between the crescent and circular shapes observed during the intermittent motion. We introduced the quantities that allow us to measure the degree of deformation and evaluated them by image processing of the experimental results. Then, we discussed the relationship between these quantities and the droplet speed. We also measured the interfacial tension between water and paraffin containing oil red O and proposed the deformation mechanism towards a crescent shape coupled with intermittent motion.
@@ II.1 Spontaneous motion of a droplet
(+)-Camphor and oil red O (dye) were purchased from FUJIFILM Wako Pure Chemical Corporation (Japan), and paraffin was purchased from Sigma-Aldrich (USA). The chemicals were used without further purification. Ion-free water was obtained using Elix UV3 (Merck, Germany). We prepared a paraffin solutions of camphor and oil red O at different concentrations by stirring using a vortex mixer and sonication.
We observed the time evolution of $20~{}\mu\mathrm{L}$ droplets located on the water surface in a Petri dish with an inner diameter of $194.5~{}\mathrm{mm}$. The time evolution of the system was recorded from above for an hour at $50~{}\mathrm{fps}$ with the CMOS camera (STC-MBS43U3V, OMRON SENTECH Co., LTD., Japan) equipped with the objective lens (3Z4S-LE SV-0814H, OMRON Co., LTD., Japan). The Petri dish was illuminated by a light panel placed below.
All experiments were performed at room temperature ($22\pm 2{}^{\circ}\mathrm{C}$). The recorded videos were analyzed using an image-processing software ImageJ[43].
@@ II.2 Interfacial tension between oil and water
The interfacial tension between pure water and oil was measured using the pendant drop method. Paraffin containing oil red O was ejected through a hook-shaped needle (4983 PD 22G, Kyowa Interface Science Co., Japan) into the aqueous phase. The droplet image was recorded from the side with the above-mentioned camera equipped with the objective lens (TEC-M55, CBC Co., Ltd., Japan). It was illuminated by the LED lamp (AS3000, AS ONE Co., Japan) located at the opposite side to the camera. The interfacial tension was calculated by fitting the shape of an oil droplet to a theoretical curve. The details are described in Appendix A.
@@ III.1 Spontaneous motion of a droplet
A $20~{}\mu\mathrm{L}$ droplet of paraffin with dissolved camphor (concentration $0.06~{}\mathrm{M}$) and oil red O (concentration $1.0~{}\mathrm{g/L}$) exhibited three different types of motion depending on the time $t$ elapsed from the moment when the droplet was placed on the water surface. At the beginning of the experiment, the droplet showed self-propelled motion along a straight line perturbed by reflections with the dish wall, as illustrated in Fig. 1(a-1). Approximately at $t=30~{}\mathrm{s}$, the droplet began to rotate along the wall (Fig. 1(a-2)). Then, at $t\sim 2000~{}\mathrm{s}$, the droplet motion switched to intermittent one (Fig. 1(a-3)). At this stage, short intervals of time when a droplet moved (less than $10~{}\mathrm{s}$) were separated by periods when it was standing. The droplet had a circular shape when it did not move, and it became elongated and exhibited deformation into a crescent shape during intermittent motion.
@@
Figure 1: Typical long-time behavior of a paraffin droplet containing camphor (concentration $0.06~{}\mathrm{M}$) and oil red O (concentration $1.0~{}\mathrm{g/L}$) moving on a water surface. The time $t=0$ corresponds to the moment when the droplet was placed on the water surface. (a) Snapshots of the droplet motion at around $t\sim 0~{}\mathrm{s}$ (a-1), $t\sim 600~{}\mathrm{s}$ (a-2), and $t\sim 2100~{}\mathrm{s}$ (a-3). The superposed images of the droplet are shown with the time difference of $0.5~{}\mathrm{s}$ (a-1) and $1.0~{}\mathrm{s}$ (a-2, a-3). The red lines indicate the trajectories of the droplet center of mass (COM) position. (b) Droplet speed $v$ as a function of time $t$.
@@
Hereafter, we focus on the intermittent motion shown in Fig. 1(a-3). We defined a single phase of intermittent motion based on the droplet speed as a function of time (cf. Fig. 3(b)). If the droplet stopped ($v\leq 2~{}\mathrm{mm/s}$) for more than $4~{}\mathrm{s}$, then we assumed that the droplet was at the rest state. If the droplet at the rest state accelerated and its speed exceeded $10~{}\mathrm{mm/s}$, then we considered this moment as the start of the intermittent phase. In the case that the time duration between the start of motion and the return to the rest state was less than $10~{}\mathrm{s}$, we defined that the droplet had performed an intermittent motion, and we analyzed the time series during each intermittent motion in detail. To minimize the boundary effect we did not analyze the intermittent motion if the droplet got closer than $20~{}\mathrm{mm}$ to the wall of the Petri dish. In case that a droplet exhibited fission, we excluded the data after the fission.
@@
In the following analysis, we focused on the correlations between the droplet speed and dynamic deformation into a crescent shape observed during the intermittent motion (Fig. 1(a-3)). The observed droplet elongation was always perpendicular to the direction of motion. For quantitative study on the deformation dynamics, we introduced the local coordinates associated with the direction of motion. Each frame of the observed time evolution was processed so that pixels representing the droplet were black (the logical value $1$) and all the others were white (the logical value 00). For each frame, we calculated the droplet center $\bm{r}_{\mathrm{c}}(t)$ as the center of mass (COM) of the black pixels. The set of droplet centers defined the droplet trajectory. The origin of the local coordinate system was set at $\bm{r}_{\mathrm{c}}(t)$. The $Y$-axis was defined as parallel to a fitting line to the droplet trajectory with a least-square method and directed as the net displacement vector. The $X$-axis was set perpendicular to the $Y$-axis. Using these coordinates, we defined the time-dependent aspect ratio $R=a/b$, where $b$ is the width of the droplet measured on the $Y$-axis and $a$ is the length of the droplet projected on the $X$-axis as illustrated in Fig. 2(a). The second quantity describing time-dependent droplet shape is the front-back asymmetric parameter $A=\Delta\Sigma/(2\Sigma)$, where $\Sigma$ is the area (the number of black pixels) of the droplet image and $\Delta\Sigma$ is the area of the figure obtained by applying XOR operation to the original droplet image and its image reflected with respect to the $X$-axis as shown in Fig. 2(b). It should be noted that $0\leq A\leq 1$ and $A=0$ for a droplet with mirror symmetry on the $X$-axis.
@@
Figure 3: Dynamics of a single phase of the intermittent motion of a $20~{}\mu\mathrm{L}$ paraffin droplet with camphor and oil red O. The concentrations of camphor and oil red O were $0.06~{}\mathrm{M}$ and $1.0~{}\mathrm{g/L}$, respectively. (a) The speed $v$ of the droplet COM, the aspect ratio $R$, and the asymmetric parameter $A$ as functions of $\Delta t$. The time $\Delta t=0$ corresponds to when $v$ has the maximum value. (b) A sequence of frames showing droplet location and deformation during the phase of intermittent motion described in (a).
@@
Figure 4: Scatter plot of $\Delta t_{A}$ and $\Delta t_{R}$, which are times for maximum values of $A(t)$ and $R(t)$ for the paraffin droplet with camphor and oil red O. The camphor concentration was $0.06~{}\mathrm{M}$ and oil red O concentration was $1.0~{}\mathrm{g/L}$.
Next, we investigated the dependence of the droplet deformation on the oil red O concentration. We performed experiments at oil red O concentrations in the range of 0.50.50.50.51.$1.1~{}\mathrm{g/L}$. The maximum values of $R(t)$ and $A(t)$ ($R_{\mathrm{max}}$ and $A_{\mathrm{max}}$), increased for oil red O concentrations larger than $0.8~{}\mathrm{g/L}$ as shown in Fig. 5.
Figure 5: (a) Maximum values of $R_{\mathrm{max}}$ and (b) $A_{\mathrm{max}}$, for $R(t)$ and $A(t)$ as functions of oil red O concentration. The camphor concentration was $0.06~{}\mathrm{M}$, and experiments at each oil red O concentration were repeated more than $5$ times. The error bars indicate the standard deviation of the results.
@@ III.2 Interfacial tension between oil and water
To understand the role of oil red O during the droplet evolution, we measured the interfacial tension $\gamma_{\mathrm{ow}}(t)$ between a paraffin droplet containing oil red O and pure water using the pendant drop method[44] as a function of time. The method is described in Appendix A. The measurements were performed for various oil red O concentrations. For the oil red O concentration $\geq 0.5~{}\mathrm{g/L}$, the interfacial tension gradually decreased in time after the interface was formed as shown in Fig. 6. Assuming that the diffusional transport of oil red O molecules and the linear absorption process of them at the droplet interface are responsible for the decrease in the interfacial tension, the time evolution of the interfacial tension should be described as an exponential relaxation. Thus, for quantitative evaluation, we fitted time-dependent interfacial tension $\gamma_{\mathrm{ow}}(t)$ to an exponential function
@@
where $\gamma_{\mathrm{ow}}^{(0)}$ is the interfacial tension at the initial stage, $\Delta\gamma_{\mathrm{ow}}$ is the difference in interfacial tension between the initial and final stages, and $\tau$ is the characteristic time for the decrease in interfacial tension. We plotted $\gamma_{\mathrm{ow}}^{(0)}$, $\Delta\gamma_{\mathrm{ow}}$ and $\tau$ against oil red O concentration in Fig. 7. $\gamma_{\mathrm{ow}}^{(0)}$ and $\Delta\gamma_{\mathrm{ow}}$ were decreasing and increasing functions of oil red O concentration, respectively. In the case when the decrease in interfacial tension was clearly observed (oil red O concentrations: $0.5$, $0.8$, and $1.0~{}\mathrm{g/L}$), $\tau$ was approximately equal to $800~{}\mathrm{s}$. In addition, we performed the same measurement for the oil droplet that contained both oil red O ($1.0~{}\mathrm{g/L}$) and camphor ($0.06~{}\mathrm{M}$). The decrease in interfacial tension was also observed, and $\gamma_{\mathrm{ow}}^{(0)}$ was lower compared to paraffin droplet with oil red O ($1.0~{}\mathrm{g/L}$) but without camphor.
Figure 6: Time evolution of the interfacial tension $\gamma_{\mathrm{ow}}(t)$ between a paraffin droplet and pure water measured by the pendant drop method. The time $t=0$ corresponds to when a droplet was formed. Marker colors code the oil red O concentrations: $0~{}\mathrm{g/L}$ black, $0.1~{}\mathrm{g/L}$ yellow, $0.5~{}\mathrm{g/L}$ orange, $0.8~{}\mathrm{g/L}$ red, $1.0~{}\mathrm{g/L}$ dark red. For each concentration, the average values of four experiments are plotted. The error bars indicate the standard deviation of the results. $\gamma_{\mathrm{ow}}(t)$ between pure water and paraffin droplet that contained both oil red O ($1.0~{}\mathrm{g/L}$) and camphor ($0.06~{}\mathrm{M}$) is marked by
blue squares.
Figure 7: Parameters fitting $\gamma_{\mathrm{ow}}(t)$ according to Eq. (1) for different concentrations of oil red O (cf. Fig. 6). (a) Interfacial tension on the initial stage $\gamma_{\mathrm{ow}}^{(0)}$. (b) Difference in interfacial tension $\Delta\gamma_{\mathrm{ow}}$ between the initial and final stages. (c) Characteristic time $\tau$ for the decrease in interfacial tension. The data for a paraffin droplet that contained both oil red O ($1.0~{}\mathrm{g/L}$) and camphor ($0.06~{}\mathrm{M}$) are shown with blue squares. The average values for four measurements are shown. The error bars indicate their standard deviations.